A toy company wishes to put a new toy in the market for the Christmas season. It does not want to do so if the manufacturing process produces more than 10% of toys that do no work. In a test run producing 80 of the toys, 10 are found to be defective. Based on the sample evidence, should the company continue to put a new toy in the market? Use 0.05 level of significance.


1
Expert's answer
2022-06-15T16:28:42-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\le0.1"

"H_a:p>0.1"

This corresponds to a right-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a right-tailed test is "z_c = 1.6449."

The rejection region for this right-tailed test is "R = \{z: z>1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{10/80-0.1}{\sqrt{\dfrac{0.1(1-0.1)}{80}}}=2.3570"

Since it is observed that "z=2.3570> z_c = 1.6449," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(Z>2.3570)=0.009212," and since "p =0.009212< 0.05," it is concluded that the null hypothesis is rejected.


It is concluded that the null hypothesis Ho is rejected.

Therefore, there is enough evidence to claim that the population proportion "p" is greaterthan "0.1," at the "\alpha = 0.05" significance level.